OpenStudy (anonymous):

can anyone exsplain how you would use the elimination method to solve these equations 1) 5x + 2y = 5 & 2) 3x - 4y = -23

OpenStudy (radar):

Which variable do you wish to eliminate, x or y? Your choice.

OpenStudy (radar):

I personally would choose y, but you may prefer x. Which one.....choose so we can start!

OpenStudy (anonymous):

lets do Y!

OpenStudy (radar):

Good choice, but x would of been ok also. Y it is, we must get the coefficient of y on one equation to equal the coefficient of y in the other equation. We must ultimately have equal coefficients for y.

OpenStudy (radar):

To do that, we will multiply the first equation by 2. Getting:10x+4y=10 Note that we now have 4 for the coefficient of y in both equations. Do you follow with understanding?

OpenStudy (anonymous):

i do

OpenStudy (radar):

10x+4y = 10 3x - 4y =-23

OpenStudy (radar):

These are the two equations that we will use to eliminate Y All we have to do is to add those two equations and the Y will just disappear (have a 0 coefficient)

OpenStudy (anonymous):

alright!

OpenStudy (radar):

Adding them I get 13x=-13 so x = -1

OpenStudy (radar):

Now substitute the -1 for x on one of the original equations and solve for Y. Now ur good.

OpenStudy (anonymous):

so id take x and use -1 on 10x ? like this 10 -1 = 9?

OpenStudy (radar):

I would take the original equation 5x+2y=5 (you could of used the other one just as well) Now substisute the value found for x (the -1) getting: 5(-1) + 2y =5 using algebra to solve: -5 +2y = 5 2y=10 y=5 the answer is y=5, it should work also with the second equation.

OpenStudy (radar):

Just to confirm. 3x-4y=-23 3(-1)-4y=-23 -3-4y=-23 -4y=-20 y=5 as you see it works, it is a good answer.

OpenStudy (anonymous):

wow thanks a ton man. helped alot.

OpenStudy (radar):

Summarizing: 1. Select a variable to eliminate. 2. Multiply or divide the equation to obtain equal coefficients (if they are opposite sign OK) 3. Add or subtract the two equations to eliminate the coefficients (if sign are opposite add, if they are same subtract) 4. Solve for remaining variable 5. substitute new found value into original equation and solve for other variable.

OpenStudy (radar):

@Luis Rivera what is that about???